Writers' Company produces 2 products presently
Fountain
Pen ball point
Unit 10000 20000
Area occupied (sq feet) 7000 8000
Variable cost - direct 32 25
Fixed cost - direct (per unit) 10 8

Rent – fixed cost per
Month(indirect – common for all products)

Compute cost per unit of each product using Activity Based Costing

Break-up of Indirect Fixed Costs Allocation basis

Rent 3,00,000 Area of operations
1
Expert's answer
2020-12-15T07:13:35-0500

Total variable cost = "variable cost per unit \times units" = "Rs 32\times 10000 units" = "Rs 320,000"

Total fixed cost = "fixed cost per unit \times units = Rs 10 \times 10000 units = Rs 100,000"

Total Fountain’s rent = 

Total products’ units = 10000 + 20000 = 30000

Total Fountain’s rent = "\frac{10000}{30000} \times Rs 300,000 = Rs 100,000"

Total cost = total variable cost + total fixed cost + total rent

= Rs 320,000 + Rs 100,000 + Rs 100,000

= Rs 520,000

Cost per unit = "\frac{Total cost}{total units} = \frac{ Rs 520,000}{10000 units } =Rs 52"


Pen ball point’s cost per unit

Total variable cost = "variable cost per unit \\times \tunits = Rs 25 \\times 20000 units = Rs 500,000"

"Total fixed cost = fixed cost per unit \times units = Rs 8 \times 20000 units = Rs 160,000"

Total Pen ball point’s rent = 

"\frac{ Pen ball point’s units }{ total products’ units } \times the total fixed indirect cost of rent"

Total products’ units = 10000 + 20000 = 30000

Total Pen ball point’s rent = "\frac{ 20000}{ 30000} \times Rs 300,000 = Rs 200,000"

Total cost = total variable cost + total fixed cost + total rent

= Rs 500,000 + Rs 160,000 + Rs 200,000

= Rs 860,000

"Cost per unit = \frac{Total cost }{ total units } = \frac{ Rs 860,000}{ 20000 units } = Rs 43"